tickets for a play cost $2 for students and $5 for adults. if a total of 500 tickets are sold and the generated revenue is $1300, how many student and adult tickets are sold.
Answer by TimothyLamb(4379)

2s + 5a = 1300
s + a = 500
s = 500 - a
---
2(500 - a) + 5a = 1300
1000 - 2a + 5a = 1300
1000 + 3a = 1300
3a = 300
Answer:
a = 100
s = 400
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php